OpenStudy (anonymous):

Find the distance between the points (1, 3) and (-7, 2)

OpenStudy (anonymous):

\[\sqrt{65}\] \[\sqrt{37}\] \[\sqrt{17}\]

OpenStudy (amistre64):

subtract one point from the other, then square add sqrt

OpenStudy (amistre64):

(1, 3) -(-7, 2) --------- 8,1; 8^2 + 1^2 = 65 sqrt(65)

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

the method works for any 2 points, so feel free to apply it :)

OpenStudy (anonymous):

can u answer mine? in a given triangle, the point of intersection of the three medians is the same as the point of intersection of three altitudes in which of the following triangles. scalene, isosceles, equilateral, or right isosceles?

OpenStudy (amistre64):

lets move it to your questions post tho

OpenStudy (anonymous):

thanks. i have a few question if you dont mind helping me. thanks.